OpenStudy (anonymous):

Let f be a continuous function such that f(1) = 3/2 and f(1/2^0.5) = 3. Then the value of \[\lim_{x \rightarrow 0}f(\frac{ \sqrt{x+1}-\sqrt{1-x} }{x })\]

OpenStudy (anonymous):

@hartnn

OpenStudy (dape):

Since f is continuous, we must have that \[ \lim_{x\rightarrow0}f(\frac{\sqrt{x+1}-\sqrt{1-x}}{x})=f(\lim_{x\rightarrow0}\frac{\sqrt{x+1}-\sqrt{1-x}}{x}) \] Applying L'Hopitals rule to the limit, \[ \lim_{x\rightarrow0}\frac{\sqrt{x+1}-\sqrt{1-x}}{x}=\lim_{x\rightarrow0}\frac{\frac{1}{2\sqrt{x+1}}-\frac{1}{2\sqrt{1-x}}}{1}=\frac{1}{2\sqrt{0+1}}-\frac{1}{2\sqrt{1-0}}=1 \] So the initial limit simply is \[ f(1)=\frac{3}{2} \]

OpenStudy (anonymous):

Exactly